# The Impact of Geometric Complexity on Neural Collapse in Transfer Learning

### v1

@@ 2 Background and Related Work
Neural collapse [22, 45, 27] refers to a phenomenon observed in deep learning where the embedding network (i.e., the subnetwork $f$ before the logit layer $g$ in a neural network $h(x)=\operatorname{softmax}(g(f(x))$) collapses the input points around their repective class means. Furthermore, these class means form a simplex creating an equiangular tight frame (ETF) centered around the global mean with roughly equal distance between its vertices. Intuitively, such a phenomenon is beneficial to generalization since the embeddings of different classes are optimally separated, making the job of the classifier head $\operatorname{softmax}(g(z))$ easier. To measure neural collapse, the authors in [22] introduce the class-distance normalized variance ($\operatorname{CDNV}$) as
$V_{f}(Q_{i},Q_{j}):=\frac{\operatorname{Var}_{f}(Q_{i})+\operatorname{Var}_{f}(Q_{j})}{\|\mu_{f}(Q_{i})-\mu_{f}(Q_{j})\|^{2}},$ (1)
@@
$\operatorname{GC}(f,Q):=\mathbb{E}_{x\sim Q}\left[\|\nabla_{x}f(x)\|^{2}\right].$
Intuitively, the GC measures the function complexity or variability and is closely related to the Dirichlet energy in geometric analysis [21]. Provided mild Lipschitz smoothness assumptions [41, Proposition 3.4], the GC can be estimated accurately on batches of training data $D$ by its empirical counterpart
$\widehat{\operatorname{GC}}(f,D)=\frac{1}{|D|}\sum_{x\in D}\|\nabla_{x}f(x)\|^{2},\quad\text{where }D\text{ is an i.i.d.~{}sample drawn from }Q.$ (4)
@@
In fact, the Poincaré inequality plays a central role in the proof of the generalization bound and we also make use of this property in the current work. For completeness, we recall it here, cf. [19]. A distribution $Q$ satisfies a Poincaré inequality if, for all differentiable functions $v$ defined on $\operatorname{supp}(Q)$, there exists a constant $c$ such that
$\operatorname{Var}_{v}(Q)\leq c\mathbb{E}_{x\sim Q}\left[\|\nabla_{x}v\|^{2}\right]=c\operatorname{GC}(v,Q).$ (5)
@@ 3 Problem Formulation
For the first stage of our inquiry (Section 4), we are concerned with a $k$-classification task. Let $D:=\{(x_{i},y_{i})\}_{i=1}^{m}$ be a dataset drawn from a distribution $Q$ with $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{e_{j}~{}|~{}j=1,\dots,k\}$ where $e_{j}\in\mathbb{R}^{k}$ are the canonical basis vectors representing a one-hot encoding of the labels. We aim to learn this task using a neural network denoted by a function $h_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ parameterized by $\theta$ (though to simplify notation we subsequently drop this dependence on $\theta$). We can write $h_{\theta}(x)=\operatorname{softmax}(g(f(x))$ where $g:\mathbb{R}^{p}\to\mathbb{R}^{k}$ is a classifier head mapping from the feature space $\mathbb{R}^{p}$ to the prediction layer $\mathbb{R}^{k}$ and $f:\mathbb{R}^{d}\to\mathbb{R}^{p}$ is a feature mapping from the input space to the penultimate embedding layer of the network before the classification head. The standard way of training $h_{\theta}$ is to find, via some stochastic optimization technique, an optimal parameter configuration $\theta_{*}$ such that $\theta_{*}=\operatorname{argmin}_{\theta}\sum_{i=1}^{m}\ell(h_{\theta}(x_{i}),y_{i})$ for a given loss function $\ell:\mathbb{R}^{k}\times\mathbb{R}^{k}\to[0,\infty)$.
@@ 4.1 Geometric complexity controls neural collapse
$\frac{\operatorname{GC}(f,Q)}{k-1}\left(\sum_{i\neq j}\frac{1}{d_{ij}^{2}}\right).$ (7)
@@
We verify the relationship posed in Eq. (6) through experiments on VGG-13 trained on CIFAR-10 (see Figure 1). Through both implicit and explicit regularizers the pressure on the embedding $\operatorname{GC}$ translates to a direct pressure on the neural collapse of the model via the geometric collapse. In Appendix A.5 we see the same relationship holds across various datasets; e.g., MNIST, Fashion-MNIST, CIFAR-100, and architectures; e.g., ResNet-18, ResNet-50.
@@ 4.2 The geometric complexity is a reliable and robust measure
One of the key advantages of the GC as a complexity measure is its computational efficiency. Under mild regularity constraints on the model function the theoretical geometric complexity of a map can be efficiently estimated by its empirical counterpart, as stated in the proposition below, already proven in [41] for logit networks (see Appendix A.2 for a proof). Additionally and crucially, we verify that the empirical $\operatorname{GC}$ is a robust and reliable measure with respect to that data sample.
@@ 4.3 A new generalization bound with GC through NC
$\mathbb{E}\left[\operatorname{Error}\right]\leq 16c\left(\frac{1}{p}+\frac{1}{m_{c}}\right)\left(\widehat{\operatorname{GC}}(f,S)+L\sqrt{\frac{\log\frac{1}{\delta}}{2km_{c}}}\right)\left(\sum_{i\neq j}\frac{1}{d_{ij}^{2}}\right),$ (10)
where $p$ is the embedding dimension of the feature map $f$ and $d_{ij}=\|\mu_{f}(S_{i})-\mu_{f}(S_{j})\|$.
@@ 5.1 Lower pre-trained GC leads to improved fine-tuning
However these other two terms involve both the source and target classes, making the full bound no longer dependent only on the geometric complexity over $\tilde{Q}$. Moreover these two terms can apriori explode since the infimum over the class-mean distances $\Delta(\mathcal{F}^{*})$ no longer only involves source classes (where this distance is bounded away from zero due to neural collapse into an equidistant simplex) but also target classes (for which we no longer have theoretical guarantees).
@@ 5.2 Improve fine-tuning by controlling pre-trained GC
6 Conclusion
Uncovering and understanding the implicit biases that enable successful transfer learning is a critical area of research in modern machine learning. Here we provide a framework connecting three different themes behind implicit regularization: flatness of the loss surface, geometric complexity, and neural collapse. We show that the embedding geometric complexity directly controls the neural collapse during training and, moreover, plays a role in the success of transfer learning. Extensive experiments on different image classification and fine tuning tasks across different hyperparameters verify our hypothesis. We believe this opens up an intriguing direction of research further exploring the role of geometric collapse and geometric complexity in deep learning and provides valuable insight for designing more efficient and effective techniques for model training and fine-tuning.
@@ Appendix A Appendix
$\operatorname{Var}_{f}(Q)\leq c\mathbb{E}_{Q}(\|\nabla_{x}f\|^{2})=c\operatorname{GC}(f,Q)$
@@ A.4 Experiment details
A.4.2 Figure 3: Controlling target performance through source GC on CIFAR-FS.

### v3

@@ 2 Background and Related Work
Neural collapse [22, 46, 27] refers to a phenomenon observed in deep learning where the embedding network (i.e., the subnetwork $f$ before the logit layer $g$ in a neural network $h(x)=\operatorname{softmax}(g(f(x))$) collapses the input points around their respective class means. Furthermore, these class means form a simplex creating an equiangular tight frame (ETF) centered around the global mean with roughly equal distance between its vertices. Intuitively, such a phenomenon is beneficial to generalization since the embeddings of different classes are optimally separated, making the job of the classifier head $\operatorname{softmax}(g(z))$ easier. To measure neural collapse, the authors in [22] introduce the class-distance normalized variance ($\operatorname{CDNV}$) as
$V_{f}(Q_{i},Q_{j}):=\frac{\operatorname{Var}_{f}(Q_{i})+\operatorname{Var}_{f}(Q_{j})}{2\|\mu_{f}(Q_{i})-\mu_{f}(Q_{j})\|^{2}},$ (1)
@@
$\operatorname{GC}(f,Q):=\mathbb{E}_{x\sim Q}\left[\|\nabla_{x}f(x)\|_{F}^{2}\right].$
Intuitively, the GC measures the function complexity or variability and is closely related to the Dirichlet energy in geometric analysis [21]. Provided mild Lipschitz smoothness assumptions [42, Proposition 3.4], the GC can be estimated accurately on batches of training data $D$ by its empirical counterpart
$\widehat{\operatorname{GC}}(f,D)=\frac{1}{|D|}\sum_{x\in D}\|\nabla_{x}f(x)\|_{F}^{2},\quad\text{where }D\text{ is an i.i.d.~{}sample drawn from }Q.$ (4)
@@
In fact, both this work and the proof of the generalization bound rely on the Poincaré inequality, which we argue is indeed a natural and mild assumption for the types of data distributions typically encountered in machine learning; see Appendix A.6 for further discussion. For completeness, we recall it here, cf. [19]. A distribution $Q$ satisfies a Poincaré inequality if, for all differentiable functions $v$ defined on $\operatorname{supp}(Q)$, there exists a constant $c$ such that
$\operatorname{Var}_{v}(Q)\leq c\mathbb{E}_{x\sim Q}\left[\|\nabla_{x}v\|_{F}^{2}\right]=c\operatorname{GC}(v,Q).$ (5)
@@ 3 Problem Formulation
For the first stage of our inquiry (Section 4), we are concerned with a $k$-classification task. Let $D:=\{(x_{i},y_{i})\}_{i=1}^{m}$ be a dataset drawn from a distribution $Q$ with $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{e_{j}~{}|~{}j=1,\dots,k\}$ where $e_{j}\in\mathbb{R}^{k}$ are the canonical basis vectors representing a one-hot encoding of the labels. We aim to learn this task using a neural network denoted by a function $h_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ parameterized by $\theta$ (though to simplify notation we subsequently drop this dependence on $\theta$).
We can write $h_{\theta}(x)=\operatorname{softmax}(g(f(x))$ where $g:\mathbb{R}^{p}\to\mathbb{R}^{k}$ is a classifier head mapping from the feature space $\mathbb{R}^{p}$ to the prediction layer $\mathbb{R}^{k}$ and $f:\mathbb{R}^{d}\to\mathbb{R}^{p}$ is a feature mapping from the input space to the penultimate embedding layer of the network before the classification head. The standard way of training $h_{\theta}$ is to find, via some stochastic optimization technique, an optimal parameter configuration $\theta_{*}$ such that $\theta_{*}=\operatorname{argmin}_{\theta}\sum_{i=1}^{m}\ell(h_{\theta}(x_{i}),y_{i})$ for a given loss function $\ell:\mathbb{R}^{k}\times\mathbb{R}^{k}\to[0,\infty)$.
@@ 4.1 Geometric complexity controls neural collapse
$\frac{\operatorname{GC}(f,Q)}{k-1}\left(\sum_{i\neq j}\frac{1}{d_{ij}^{2}}\right),$ (7)
where $k$ is the number of classes, and $d_{ij}=\|\mu_{f}(Q_{i})-\mu_{f}(Q_{j})\|$ denotes the Euclidean distance between the mean $\mu_{f}(Q_{i})$ and $\mu_{f}(Q_{j})$ of class $i$ and class $j$ (resp.) in embedding space.
@@
We verify the relationship posed in Eq. (6) through experiments on VGG-13 trained on CIFAR-10 (see Figure 1). Through both implicit and explicit regularizers the pressure on the embedding $\operatorname{GC}$ translates to a direct pressure on the neural collapse of the model via the geometric collapse. As already observed in [17], lower levels of GC coincide with higher test accuracy as show in Figure 5 in Appendix A.5 containing the learning curves of that experiment.
In Appendix A.5 we see the same relationship holds across various datasets; e.g., MNIST, Fashion-MNIST, CIFAR-100, and architectures; e.g., ResNet-18, ResNet-50. To avoid possible confounding factors that could arise through batch size and learning rate manipulations, we also directly regularize with the geometric complexity in Appendix A.5.6; the same relations hold in this case too.
@@ 4.2 The geometric complexity is a reliable and robust measure
One of the key advantages of the GC as a complexity measure is its computational efficiency.
Under mild regularity constraints on the model function the theoretical geometric complexity of a map can be efficiently estimated by its empirical counterpart, as stated in the proposition below, already proven in [42] for logit networks (see Appendix A.2 for a proof). Additionally and crucially, we verify that the empirical $\operatorname{GC}$ is a robust and reliable measure with respect to that data sample.
@@ 4.3 A new generalization bound with GC through NC
$\mathbb{E}\left[\operatorname{Error}\right]\leq 16c\left(\frac{1}{p}+\frac{1}{m_{c}}\right)\left(\widehat{\operatorname{GC}}(f,S)+L\sqrt{\frac{\log\frac{2}{\delta}}{2m_{c}k}}\right)\left(\sum_{i\neq j}\frac{1}{d_{ij}^{2}}\right),$ (10)
where $p$ is the embedding dimension of the feature map $f$ and $d_{ij}=\|\mu_{f}(S_{i})-\mu_{f}(S_{j})\|$. Note, the outer expectation on the left hand side is taken over the samples $S$ used to produce the classifier $h_{f,S}$.
@@
In Figure 3, we plot the LHS and RHS of the generalization bound in (10) from Proposition 4.3 for a VGG-13 model trained on CIFAR-10. In this plot, we see that the bound is not vacuous, demonstrating a relatively tight fit. Note that on the RHS, we omitted the term involving the Lipschitz constant $L$, assuming it is small due to its denominator. Additionally, we estimated a lower bound for the Poincaré constant $c$ by comparing the magnitudes of the RHS and LHS in the inequality (6), based on results in Figure 1. This approximation yielded $c\approx 1000$ in our setup.
Figure 3: VGG-13 trained on CIFAR-10 with 5 random seeds.
@@ 5.1 Lower pre-trained GC leads to improved fine-tuning
However these other two terms involve both the source and target classes, making the full bound no longer dependent only on the geometric complexity over $\tilde{Q}$.
Moreover these two terms can apriori explode since the infimum over the class-mean distances $\Delta(\mathcal{F}^{*})$ no longer only involves source classes (where this distance is bounded away from zero due to neural collapse into an equidistant simplex) but also target classes (for which we no longer have theoretical guarantees).
@@ 5.2 Improve fine-tuning by controlling pre-trained GC
6 Limitations and Conclusion
There is a notable limitation when extending our findings to language modeling and, for example, large language models (LLMs). However, this limitation is not specific to our work per-se, but instead it is a limitation of the application of neural collapse to language models in general, as described in the recent work [65]. Namely, language modeling, as conducted via training by token prediction, creates a classification task where the conditions for neural collapse are implausible. The main problem, in addition to an imbalanced vocabulary, is that the embedding dimension for language models is typically far less than the number of classes (i.e., the total vocabulary size), making the neural collapse simplex impossible to exist. Extending the notion of neural collapse to the language models is an open question and an active area of research and beyond the scope of this work.
Uncovering and understanding the implicit biases that enable successful transfer learning is a critical area of research in modern machine learning. Here we provide a framework connecting three different themes behind implicit regularization: flatness of the loss surface, geometric complexity, and neural collapse. We show that the embedding geometric complexity directly controls the neural collapse during training and, moreover, plays a role in the success of transfer learning.
Extensive experiments on different image classification and fine tuning tasks across different hyperparameters verify our hypothesis. We believe this opens up an intriguing direction of research further exploring the role of geometric collapse and geometric complexity in deep learning and provides valuable insight for designing more efficient and effective techniques for model training and fine-tuning.
Acknowledgments and Disclosure of Funding
We would like to thank Patrick Cole for their support. We would also like to thank Hanna Mazzawi and Peter Bartlett for helpful conversations.
@@ Appendix A Appendix
$\operatorname{Var}_{f}(Q)\leq c\mathbb{E}_{Q}(\|\nabla_{x}f\|_{F}^{2})=c\operatorname{GC}(f,Q)$
@@ A.4 Experiment details
In Figure 5, we show the complete learning curves of that experiment.
Figure 5: VGG-13 on CIFAR-10. Training has reached terminal phase of training (TPT) with training accuracy equal to 1 (see first and second column). Lower GC in training correlates with higher test accuracy in all settings (see third and fourth column).
A.4.2 Figure 3: Tightness of the generalization bound on CIFAR-10
We trained a VGG-13 neural network [56] on the full CIFAR-10 dataset using the provided train/test split. For this model architecture we use an embedding layer dimension $p=1024$. Note also that the number of examples per class is $m_{c}=5000$. We used a constant learning rate of $0.005$, a batch size of $512$ and trained for $100000$ steps reporting metrics every $1000$ steps.
Throughout training, we reported the following metrics measured and averaged over multiple batches of the training dataset: 1) the empirical geometric complexity of the model embedding layer; i.e., the layer on which the neural collapse is measured, denoted $\widehat{\operatorname{GC}}$, and 2) the sum of the inverse squares of $d_{ij}$ which denotes the distance between the mean of class $i$ and class $j$ for $i,j\in[k]$ for $i\neq j$ and, here, $k=10$. These quantities were used to create the blue curve in Figure 3. We also computed the nearest mean classifier error on the test set. Recall, for the feature map $f$ and a sample $S$ the nearest mean classifier is defined as $h_{f,S}:=\operatorname{argmin}_{c\in[k]}\|f(x)-\mu_{f}(S_{c})\|$ where, here $k=10$, and $\mu_{f}(S_{c})$ denotes the sample mean of the set $S_{c}$ for class $c$ under the feature map $f$. To create the orange curve in Figure 3 we plot the average error $h_{f,S}(x)\neq y$ over the test set.
A.4.3 Figure 4: Controlling target performance through source GC on CIFAR-FS.